# Binary Flow Matching: Prediction-Loss Space Alignment for Robust Learning

### v1

@@ V-B Results and Discussion
As shown in Fig. 2(a–d), uniform $t$-sampling leads to severe gradient explosion and unstable loss behavior under $x$-prediction with velocity loss on both Gaussian and binary manifolds, while Logit-Normal sampling yields bounded gradients and smooth loss decay.
However, Fig. 2(e) reveals that this stabilization is achieved by heavily suppressing samples near $t\to 1$, which is precisely where the theoretical singularity occurs.
Furthermore, Fig. 2(f–i) confirms that aligned prediction–loss configurations remain stable even under uniform sampling, indicating that the instability is not inherent to the data distribution or network architecture, but arises from structural mismatch. Finally, Fig. 2(j) shows that performance degradation in the binary case concentrates near the boundary region, consistent with our analysis of residual-driven singularity.
These results support the conclusion that Logit-Normal sampling mitigates instability by avoiding the singular region, rather than resolving the underlying mismatch, motivating the sampler-agnostic stabilization provided by prediction–loss space alignment.
VI Experiments
@@
In this section, we evaluate two representative problems: binary MNIST handwritten digit generation and multiple-input multiple-output (MIMO) signal detection. These tasks contrast the strong spatial correlations of visual data with the symbol-wise independence of communication signals.
Our results reveal that prediction-loss alignment is a fundamental prerequisite for robust learning; standard heuristics like Logit-Normal sampling or $t\to 1$ clipping merely mask singularities without resolving the underlying numerical fragility. Furthermore, we demonstrate that data topology dictates the optimal choice between BCE and MSE once alignment is secured. While not aiming for state-of-the-art generative benchmarks, these results establish alignment as a key principle for stable and topology-aware discrete flow matching.
VI-A Binary MNIST Image Generation
@@
Results and Discussion. Table I and Fig. 3 demonstrate the clear superiority of the proposed alignment principle. Our aligned $x$-prediction with $x_{\mathrm{MSE}}$ loss not only ensures monotonic convergence but also achieves the the lowest FID score among all tested configurations. This performance edge stems from the well-conditioned loss landscape, which allows the model to accurately capture fine-grained spatial structures without being hindered by numerical instability.
In stark contrast, we analyze the mismatched configuration ($x$-pred with $v_{\mathrm{MSE}}$ loss) under the Logit-Normal schedule. While the “best” checkpoint in this setting can yield a reasonably legible FID, it consistently fails to match the fidelity of our aligned approach. Furthermore, its training is characterized by extreme validation instability, with losses fluctuating across two orders of magnitude (Fig. 3, Plot 2). This suggests that even the best results from the mismatched setting are ”fragile” and highly sensitive to early-stopping.
The gap in FID underscores a key insight: our Alignment does not merely provide a “palliative masking” of the singularity; it enables the model to effectively supervise the entire probability path. By eliminating the structural singularity, our method allows for more precise learning of the deterministic transition to binary states near $t\to 1$, resulting in the sharper strokes and higher sample fidelity reflected in the superior FID. (See Appendix E for full settings.)
@@ VI-B MIMO Detection
In contrast to binary image generation, BCE-based objectives outperform MSE in MIMO detection (Fig. 4(c,d)). Since bits are i.i.d. after real-valued conversion, symbol-wise likelihood objectives better match the data distribution, while regression losses implicitly assume continuous manifolds. This further highlights that effective flow-matching on binary domains requires alignment between model design, loss function, and data topology. (See Appendix E for the SGT architecture and complete experiment settings.)
VII Conclusion
We studied flow matching and diffusion models on binary data manifolds and showed that signal-space prediction ($x$-prediction) remains effective when objectives are properly aligned. We identified that pairing $x$-prediction with velocity-matching loss induces a prediction–loss mismatch that causes gradient variance explosion near the terminal region, explaining the instability observed in practice. While heuristic sampling or timestep clipping can partially mitigate this issue, they do not resolve the underlying incompatibility. We proposed prediction-loss space alignment as a structural remedy that restores stability independently of the sampling distribution. Moreover, once alignment is enforced, the optimal loss is dictated by data topology: regression losses suit spatially correlated structures such as binary images, whereas probabilistic objectives are more appropriate for independent symbolic inference as in MIMO detection. Experiments across both tasks consistently validate these principles. Overall, our results highlight that objective design is central when extending continuous-time generative frameworks to discrete domains, and that respecting the geometric and statistical structure of binary data is essential for stable and effective learning.
@@ VII Conclusion
$\mathbb{E}[\|\mathbf{g}_{t}(\theta)\|^{2}]\geq\frac{4c}{(1-t)^{4}}R(t).$ (13)
The optimization stability is determined by the integral $\mathcal{I}=\int_{0}^{1}\mathbb{E}[\|\mathbf{g}_{t}\|^{2}]dt$.
@@
Substituting $R(t)=\mathcal{O}((1-t)^{2})$ into the integral:
@@
The integral exhibits a first-order divergence. While divergent, this instability is often numerically circumvented in practice by specialized time-sampling heuristics that avoid the $t\approx 1$ boundary.
Binary Discrete Case.
For $\mathbf{x}\in\{-1,1\}^{D}$, the network maintains a residual $R(t)\geq\epsilon^{2}$ as $t\to 1$ (Assum. IV.3). Substituting this into the integral:
$\begin{split}\mathcal{I}_{bin}&\geq\int_{t^{*}}^{1}\frac{4c\epsilon^{2}}{(1-t)^{4}}dt\\ &=\left[\frac{4c\epsilon^{2}}{3(1-t)^{3}}\right]_{t^{*}}^{1}=\infty.\end{split}$ (17)
The integral exhibits a third-order divergence. Unlike the continuous case, this explosive variance leads to immediate numerical overflow in the absence of specialized sampling or structural alignment. ∎
@@ Appendix B Asymptotic Stability Analysis of Sampling Schedules
We rigorously analyze the interaction between the Logit-Normal schedule and the mismatch singularity. Let $u=\text{logit}(t)$, then $1-t=\frac{e^{-u}}{1+e^{-u}}\sim e^{-u}$ for large $u$. The gradient variance under $\pi_{LN}(t)$ transforms as follows:
@@
For binary data, $R(u)\sim\Omega(1)$. The integrand becomes $\exp(-u^{2}/2s^{2}+4u)$. The peak is shifted further to $u_{peak}=4s^{2}$. To prevent numerical overflow from the $e^{4u}$ term, the sampler must ensure that the Gaussian tail $\exp(-u^{2})$ decays significantly before $e^{4u}$ hits the floating-point limit. This creates an Effective Sampling Gap: the model spends its ”sampling budget” in regions where $t$ is far from 1.
@@ Conclusion.
For binary flow matching, Logit-Normal sampling creates a fundamental conflict: one must choose between numerical survival (by suppressing the boundary) and binary fidelity (by sampling the boundary). Our proposed Prediction-Loss Alignment resolves this conflict by ensuring $\mathcal{O}(1)$ variance regardless of the sampling measure, enabling the model to learn the critical $t\to 1$ transition without instability.
@@ C-B Case 2: Binary Manifolds (Discrete Priors)
As in the continuous case, the residual $\|\hat{\mathbf{x}}_{\theta}-\mathbf{x}\|$ is bounded by the hypercube diameter $2\sqrt{D}$. Despite the non-vanishing residual $R(t)\sim\Omega(1)$ at the Dirac peaks, the absence of the $(1-t)^{-2}$ weighting in the objective ensures:
@@ Conclusion.
Appendix D Toy Experiment Settings
TABLE II: Toy Experiment: Network Architecture and Training Hyperparameters
@@ Appendix D Toy Experiment Settings
Appendix E Experiment Settings
Figure 5: Soft Graph Transformer Architecture
TABLE III: BMNIST flow-matching experiment configuration (Conditional UNet backbone, training, and evaluation).
@@ Appendix E Experiment Settings
TABLE IV: Hyperparameter settings for MIMO detection experiments (DiT-style SGT backbone).

### v2

@@ IV-F Controlled Verification on Toy Problems
Figure 2: Gaussian toy experiment: Logit-Normal sampling stabilizes mismatched $x$-prediction with velocity loss. (a,b) Under $x$-prediction with $v$-MSE, uniform $t$-sampling causes large gradient spikes and unstable losses, while Logit-Normal sampling yields bounded gradients and smooth convergence. (c) Logit-Normal sampling strongly suppresses the boundary region $t\to 1$, where the singular behavior is concentrated.
Figure 3: Binary toy experiment under uniform sampling: instability is specific to mismatched objectives and concentrates near the boundary. (a,b) Across prediction–loss pairings, the mismatched $x$-prediction with $v$-MSE shows much larger gradients and worse optimization, while aligned objectives remain stable. (c) BER versus $t$, together with the Bayes-optimal MMSE curve, shows that the degradation is concentrated near $t\to 1$.
On the Gaussian manifold, Fig. 2 shows that uniform $t$-sampling leads to severe gradient explosion and unstable optimization under the mismatched objective, whereas Logit-Normal sampling keeps gradients bounded. The same figure also shows why: stabilization is achieved by suppressing samples near $t\to 1$, exactly where the singularity occurs.
Fig. 3 separates sampler effects from objective mismatch. Under uniform sampling, aligned prediction–loss pairings remain stable in the binary setting, while the mismatched $x$-prediction+$v$-loss configuration alone exhibits large gradients, poor optimization, and BER degradation concentrated near the boundary. We defer the broader BCE-versus-MSE comparison to the BMNIST and MIMO experiments, where the role of data topology becomes visible. (See Appendix D-A for detailed settings.)
Figure 4: Binary MNIST qualitative samples and training dynamics under different objectives. Each column shows one prediction–loss pairing with its training/validation curves and generated samples. Mismatched objectives are unstable or brittle even under Logit-Normal sampling, whereas aligned objectives converge reliably. BCE-based models produce thicker strokes, while aligned MSE gives the best overall optimization and sample quality.
V Experiments
@@
In this section, we focus on two core binary benchmarks: Binary MNIST handwritten digit generation and multiple-input multiple-output (MIMO) signal detection. Together they contrast spatially correlated binary data and symbol-wise independent communication signals.
Our results show that prediction-loss alignment is a fundamental prerequisite for robust learning, while the preferred aligned loss still depends on data topology. We additionally include a supporting Tiny-ImageNet ablation in Appendix D-B to show that the interaction between objective mismatch and timestep sampling is also visible in a modern continuous-image diffusion architecture.
In that JiT-B/4 sanity check, the mismatched $x$-prediction+$v$-loss objective collapses under uniform timestep sampling but becomes numerically stable under Logit-Normal sampling, whereas aligned objectives remain robust. We therefore interpret Tiny-ImageNet as supporting evidence that boundary-biased sampling can mask mismatch in realistic image-generation systems, while alignment remains the more sampler-agnostic design principle. Full setup and results are deferred to Appendix D-B.
V-A Binary MNIST Image Generation
@@
Results and Discussion. We also corrected two evaluation issues from the earlier draft. First, for BMNIST there is no standard Inception-style feature extractor, so we retrained a dedicated MNIST classifier to convergence and used the best validation-accuracy checkpoint for FID computation, following common dataset-specific practice. Second, we fixed a validation bug in which the $x$-prediction+$v$-loss+Logit-Normal variant had been evaluated with uniform rather than intended validation timestep sampling. These corrections do not change the main conclusion, but they do refine some earlier comparisons among non-top variants.
[…]
Overall, the corrected BMNIST results continue to support the same main message: alignment is more stable, and once alignment is enforced, MSE is favored over BCE on strongly correlated binary manifolds. (See Appendix D-C for full settings.)
@@ V-B MIMO Detection
[…]
In contrast to binary image generation, BCE-based objectives outperform MSE in MIMO detection (Fig. 5(c,d)). Since bits are i.i.d. after real-valued conversion, symbol-wise likelihood objectives better match the data distribution, while regression losses implicitly assume continuous manifolds. This further highlights that effective flow-matching on binary domains requires alignment between model design, loss function, and data topology. (See Appendix D-C, Appendix D-C1, and Appendix D-D1 for supplementary analyses and full settings.)
VI Conclusion
We studied flow matching and diffusion models on binary data manifolds and showed that signal-space prediction ($x$-prediction) remains effective when objectives are properly aligned. We identified that pairing $x$-prediction with velocity-matching loss induces a prediction–loss mismatch that causes singular amplification near the terminal region, explaining the instability observed in practice. While non-uniform timestep sampling can mitigate this effect, our Tiny-ImageNet ablation shows that its practical benefit is configuration-dependent and should not be treated as an objective-independent improvement.
We proposed prediction-loss space alignment as a structural remedy that restores stability independently of the sampling distribution. Moreover, once alignment is enforced, the optimal loss is dictated by data topology: regression losses suit spatially correlated structures such as binary images, whereas probabilistic objectives are more appropriate for independent symbolic inference as in MIMO detection. Corrected BMNIST evaluations and fair-checkpoint MIMO comparisons continue to support these conclusions.
Overall, our results highlight that objective design is central when extending continuous-time generative frameworks to discrete domains. Our formal theory is established in the binary setting, but the additional JiT-B/4 ablation suggests that the underlying interaction between mismatch and timestep sampling is relevant more broadly in modern diffusion systems.
@@ VI Conclusion
$\mathbb{E}[\|\mathbf{g}_{t}(\theta)\|^{2}]\geq\frac{4c}{(1-t)^{4}}R_{\theta}(t),$ (13)
where $R_{\theta}(t)=\mathbb{E}\|\hat{\mathbf{x}}_{\theta}-\mathbf{x}\|^{2}$ denotes the training-stage residual. The optimization stability is determined by the integral $\mathcal{I}=\int_{0}^{1}\mathbb{E}[\|\mathbf{g}_{t}\|^{2}]dt$.
@@
Substituting $R_{\theta}(t)=\mathcal{O}((1-t)^{2})$ into the integral:
@@
The integral exhibits a first-order divergence. While divergent, this instability is often numerically circumvented in practice by specialized time-sampling heuristics that avoid the $t\approx 1$ boundary. This shows that mismatch-induced divergence persists even in the Bayes-optimal continuous benchmark.
Binary Regime A (no end-to-end residual skip).
For $\mathbf{x}\in\{-1,1\}^{D}$, assume the architecture has no explicit end-to-end skip from $\mathbf{z}_{t}$ to output, so no identity-like projection is available near initialization; then $R_{\theta}(t)\geq c_{0}>0$ on a terminal interval (Assum. IV.3). Substituting this into the integral:
$\begin{split}\mathcal{I}_{bin,A}&\geq\int_{t^{*}}^{1}\frac{4cc_{0}}{(1-t)^{4}}dt\\ &=\left[\frac{4cc_{0}}{3(1-t)^{3}}\right]_{t^{*}}^{1}=\infty.\end{split}$ (17)
The integral exhibits a third-order divergence.
Binary Regime B (with end-to-end residual skip).
If the architecture provides an explicit end-to-end residual path (e.g., identity skip with zero-initialized residual branch), then initialization gives identity-like projection $\hat{\mathbf{x}}_{\theta}\approx\mathbf{z}_{t}$, and
$R_{\theta}(t)\approx\mathbb{E}\|\mathbf{z}_{t}-\mathbf{x}\|^{2}=(1-t)^{2}\mathbb{E}\|\boldsymbol{\epsilon}-\mathbf{x}\|^{2}=\mathcal{O}((1-t)^{2}).$
Substituting into the mismatch integral:
$\begin{split}\mathcal{I}_{bin,B}&\gtrsim\int_{t^{*}}^{1}\frac{(1-t)^{2}}{(1-t)^{4}}dt=\int_{t^{*}}^{1}(1-t)^{-2}dt=\infty.\end{split}$ (18)
This corresponds to first-order divergence of $\mathcal{I}$, i.e., weaker than Regime A but still singular under mismatch. ∎
@@ Appendix B Asymptotic Stability Analysis of Sampling Schedules
We rigorously analyze the interaction between the Logit-Normal schedule and the mismatch singularity. In general, let
$u\sim\mathcal{N}(m,s^{2}),\qquad t=\sigma(u),$
so that $t$ follows a Logit-Normal distribution. In the main text we set $m=0$ for simplicity. This does not affect convergence, because $m$ only shifts the effective density in logit space and does not change the Gaussian tail responsible for suppressing the boundary singularity. Indeed,
$\exp\!\left(-\frac{(u-m)^{2}}{2s^{2}}\right)=\exp\!\left(-\frac{u^{2}}{2s^{2}}+\frac{m}{s^{2}}u-\frac{m^{2}}{2s^{2}}\right),$
so a nonzero $m$ changes only the linear and constant terms in the exponent, while the dominant negative quadratic term $u^{2}/(2s^{2})$ remains unchanged. Therefore the large-$u$ integrability is governed by $s$, not by the particular choice of $m$.
Let $u=\text{logit}(t)$, then $1-t=\frac{e^{-u}}{1+e^{-u}}\sim e^{-u}$ for large $u$. Under the simplified choice $m=0$, the gradient variance under $\pi_{LN}(t)$ transforms as follows:
@@
Binary behavior depends on training-stage residual regime. In Regime A, $R_{\theta}(u)\sim\Omega(1)$, so the integrand is $\exp(-u^{2}/2s^{2}+4u)$ with peak $u_{peak}=4s^{2}$. In Regime B, $R_{\theta}(u)\sim e^{-2u}$, so the integrand becomes $\exp(-u^{2}/2s^{2}+2u)$ with peak $u_{peak}=2s^{2}$. Thus Logit-Normal remains convergent in both regimes, while boundary suppression is most critical in Regime A.
@@ Conclusion.
For binary flow matching, Logit-Normal sampling mitigates mismatch-induced boundary amplification by suppressing large-$u$ regions, especially in Regime A. Our proposed Prediction-Loss Alignment removes the mismatch factor itself and yields $\mathcal{O}(1)$ variance regardless of the sampling measure.
@@ C-B Case 2: Binary Manifolds (Discrete Priors)
As in the continuous case, the residual $\|\hat{\mathbf{x}}_{\theta}-\mathbf{x}\|$ is bounded by the hypercube diameter $2\sqrt{D}$. Since aligned objectives do not introduce the mismatch factor $(1-t)^{-2}$, we have:
@@ Conclusion.
Appendix D Supplementary Experimental Details
D-A Toy Experiments
TABLE I: Toy Experiment: Network Architecture and Training Hyperparameters
@@ D-A Toy Experiments
Figure 6: Complete toy-experiment summary corresponding to the condensed figures in the main text. (a,b) Gradient norms under Logit-Normal versus uniform sampling for binary and Gaussian data, respectively. Uniform sampling causes severe gradient explosion under mismatched $x$-prediction with velocity loss, while Logit-Normal sampling keeps gradients bounded. (c,d) The corresponding training losses show that Logit-Normal sampling yields stable convergence, whereas uniform sampling is highly unstable. (e) Distribution of sampled $t$ values under the two timestep schedules, showing that Logit-Normal strongly suppresses the boundary region $t\to 1$. (f,g) Gradient norms under uniform sampling for different prediction–loss pairings, confirming that the instability is specific to mismatched $x$-prediction with $v$-loss. (h,i) Training losses under uniform sampling further show that aligned objectives remain stable. (j) BER as a function of $t$ in the binary case, compared with the Bayes-optimal MMSE estimator, showing that degradation is concentrated near the singular boundary region. This figure provides the full toy-experiment breakdown, while the main text keeps only the Gaussian sampling comparison and the binary mismatch/BER summary used to support the analysis.
D-B Tiny-ImageNet Ablation
To test whether the same mechanism appears beyond controlled binary settings, we additionally evaluate a JiT-B/4 Diffusion Transformer on Tiny-ImageNet ($64\times 64$). We compare three objectives under both uniform and Logit-Normal timestep sampling: mismatched $x$-prediction+$v$-loss, aligned $x$-prediction+$x$-loss, and standard flow matching $v$-prediction+$v$-loss.
Figure 7: Additional Tiny-ImageNet ablation on a JiT-B/4 backbone. We compare mismatched $x$-prediction+$v$-loss, aligned $x$-prediction+$x$-loss, and standard flow matching $v$-prediction+$v$-loss under both uniform and Logit-Normal timestep sampling. The mismatched objective collapses under uniform sampling, but becomes numerically stable under Logit-Normal sampling. In contrast, the aligned objective remains stable under both samplers and consistently outperforms standard flow matching across both sampling strategies. This highlights that timestep sampling and objective design interact in a configuration-dependent way.
Fig. 7 shows that the mismatched objective with uniform sampling collapses (FID-50K =390.50=390.50), while the same objective becomes numerically stable under Logit-Normal sampling (FID-50K =14.56=14.56). This is consistent with our theoretical claim that non-uniform timestep sampling can suppress the terminal singular region without removing the mismatch itself.
The ablation also shows that the effect of timestep sampling is configuration-dependent. Standard flow matching achieves FID-50K $23.83$ under uniform sampling but degrades to $30.43$ under Logit-Normal sampling, whereas the aligned $x$-objective achieves $15.85$ under uniform sampling and $24.32$ under Logit-Normal sampling. Thus, Logit-Normal sampling is not an objective-independent improvement; rather, it interacts non-trivially with the chosen parameterization and training loss.
We therefore view this experiment as supporting two conclusions. First, the mismatch mechanism is not specific to binary toy settings, since the same objective can collapse in a realistic image-generation system under uniform sampling. Second, alignment remains a more sampler-agnostic design principle: in this JiT-B/4 ablation it consistently outperforms standard flow matching under both timestep samplers, while the mismatched objective relies much more strongly on the Logit-Normal heuristic.
D-C Binary MNIST
TABLE II: BMNIST flow-matching experiment configuration (Conditional UNet backbone, training, and evaluation).
@@ D-C Binary MNIST
D-C1 Corrected Evaluation Protocol
For BMNIST there is no standard Inception-style feature extractor. We therefore retrain a dedicated MNIST classifier to convergence and compute FID-50K using the checkpoint with the best validation accuracy, following common dataset-specific practice. We also correct the validation bug from the earlier draft in which the $x$-prediction+$v$-loss+Logit-Normal variant had been evaluated with uniform rather than intended validation timestep sampling.
((a))
((b))
Figure 8: Corrected BMNIST evaluation protocol. We train a dedicated classifier to convergence and compute FID-50K using the best validation-accuracy checkpoint. Across classifier maturities, the aligned $x_{\mathrm{MSE}}$ objective under Logit-Normal sampling remains the strongest configuration, while some relative ordering among weaker variants changes mildly.
[…]
TABLE III: Hyperparameter settings for MIMO detection experiments (DiT-style SGT backbone).
@@ D-C1 Corrected Evaluation Protocol
D-D MIMO Detection
For completeness, we visualize both the original Soft Graph Transformer (SGT) design and the diffusion-adapted DiSGT backbone used in our MIMO experiments. The first view highlights the soft-input/soft-output message-passing structure inherited from neural MIMO detection, while the second summarizes the DiT-style conditional architecture used to implement flow matching in our setting.
Figure 9: Original Soft Graph Transformer (SGT) layer architecture. The model maintains separate soft-input and soft-output streams and updates them through encoder, cross-attention, and output blocks within each SGT layer. This view emphasizes the message-passing structure inherited from neural MIMO detection.
Figure 10: Diffusion-adapted SGT (DiSGT) architecture used in our MIMO experiments. A DiT-style encoder is combined with cross-attention to observation-dependent features and an MLP prediction head, while AdaLN modulation injects timestep and conditioning information. This condensed view highlights the architectural changes used to cast SGT as a conditional flow-matching backbone.
D-D1 Checkpoint-Fairness Analysis for $8\times 8$ MIMO
For the $8\times 8$ MIMO task, the mismatched $x$-prediction+$v$-loss configuration diverges under the main training setting, so its strongest BER is achieved at an early checkpoint. To check whether this creates a checkpoint-selection advantage for the mismatched baseline, we compare all methods both at the final training checkpoint (250K steps) and at the checkpoint that is optimal for the mismatched baseline (13K steps).
[…]
Fig. 11 shows that the qualitative conclusion is unchanged under both comparison rules. At the final checkpoint, the mismatched model fails completely; even at the 13K-step checkpoint chosen to favor that baseline, aligned objectives still maintain a clear BER advantage.